OpenStudy (anonymous):

every evening jenna empties her pocket and puts her change in a jar.at the end of the week she count her money.one week she had 38 coins,all of them dimes and quaeters.when she added them up,she had a total of $6.95 let d=number of dimes and q=number of quarters

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

man whts "dimes" and "quarters" ??

563blackghost (563blackghost):

dimes is 10 cents and quarters are 25 cents

563blackghost (563blackghost):

Is there answer choices to this question?

rishavraj (rishavraj):

100 cents = $1 ????@563blackghost

563blackghost (563blackghost):

Yes! Well in America

rishavraj (rishavraj):

@Najia2000 see it means its total 695 cents if no of dimes = x then no of quarters = 38 - x so 10x + 25(38 - x) = 695

rishavraj (rishavraj):

solve and get "x"

563blackghost (563blackghost):

dimes is to equal d quarters is to equal q

OpenStudy (texaschic101):

d + q = 38 .10d + .25q = 6.25 use substitution and solve

rishavraj (rishavraj):

@563blackghost u can solve tht way too....lol @texaschic101 its supposed to be $6.95

OpenStudy (texaschic101):

oops...my bad..... .10d + .25q = 6.95

rishavraj (rishavraj):

@Najia2000 done????

OpenStudy (anonymous):

Thank you everyone! @rishavraj i had no idea about quarters and dimes either

rishavraj (rishavraj):

hahahha lol ;P

rishavraj (rishavraj):

thanks to @563blackghost ......;P

OpenStudy (anonymous):

What can i say, aussies dont know about americans much. :p